What is the probability of correctly picking a sequence of four numbers when each number can be 0 through 9 and each is independent of the other?
Answer by stanbon(75887)

What is the probability of correctly picking a sequence of four numbers when each number can be 0 through 9 and each is independent of the other?
-------------------
# of ways to succeed 10
# of possible outcomes: 10C4 = 210
----
Probability = 10/210 = 1/21
===============================
Cheers,
Stan H.
==========